OpenStudy (anonymous):

someone explain how to simplify 9p - 2p^2 + 5p please

OpenStudy (anonymous):

Let's simplify step-by-step. 9p−2p2+5p =9p+−2p2+5p Combine Like Terms: =9p+−2p2+5p =(−2p2)+(9p+5p) =−2p2+14p

OpenStudy (anonymous):

2p2 is 2p^2

OpenStudy (anonymous):

got it. thanks :)